# FIMP Dark Matter in bulk viscous non-standard cosmologies

### v2

@@ III.1 Comparison between scenarios
Figure 7: Comparison in the yield production between the classical NSC (solid red line) and the NSC with bulk viscosity (solid blue line) for $\kappa=10^{-3}$, $T_{\text{end}}=7\times 10^{-3}$ GeV, $\hat{\xi_{0}}=10^{-2}$, $m_{\chi}=100$ GeV, $\langle\sigma v\rangle=5\times 10^{-26}$ GeV-2, and $\omega=0$. The dashed and dotted lines correspond to $x_{\text{eq}}$ (cyan), $x_{\text{c}}$ (grey), and $x_{\text{end}}$ (green) for the NSC with and without bulk viscosity, respectively; while the dashed-dotted magenta line is the time when the DM candidate Freeze-In their number at $x_{\text{fi}}$. The green zone represents the current DM relic density according to Ref. [9], observing that the parameter space considered in this case gives us the right amount of DM in the bulk viscous NSC scenario.
@@
The comparison in the parameter space of the NSC models, namely $\kappa$ and $T_{\text{end}}$, with a fixed $\omega$, is depicted in Figures 8, 9, and 10. The figures were obtained for the values $\hat{\xi_{0}}=10^{-2}$, $m_{\chi}=100$ GeV, and $\langle\sigma v\rangle=10^{-26}$ GeV-2. The solid red and blue lines correspond to the parameter space that reproduces the current DM relic density for the classical NSC and the NSC with bulk viscosity, respectively. We also delimit the three regions of interest (see Section II.1) through the equalities $T_{\text{eq}}=T_{\text{fi}}$ (cyan), $T_{\text{c}}=T_{\text{fi}}$ (grey), and $T_{\text{end}}=T_{\text{fi}}$ (green) for the NSC with bulk viscosity (dashed lines) and the classical NSC (dashed-dotted lines). The red and blue zones represent their respective parameter space in which $\rho_{\phi}<\rho_{\gamma}$, $\forall t$, i.e., the regions in which the Hubble parameter is always approximately similar to the $\Lambda$CDM model ($H\sim T^{2}$). The grey zone corresponds to the forbidden BBN epoch that starts at $T_{\text{BBN}}\sim 4\times 10^{-3}$ GeV. In Figure 8, obtained for $\omega=0$, it can be seen that for higher values of $\kappa$ and $T_{\text{end}}$ (RIII), the parameter space is slightly the same, meanwhile in RII the curves separate from each other. However, the greatest differences between the NSC with and without bulk viscosity cases are presented in RI, exhibiting an approximate independent behavior of the $\kappa$ parameter, which can be explained by the dominance of the viscous term over the decaying term. On the other hand, Figure 9, obtained for $\omega=-2/5$, exhibits a similar behavior to the $\omega=0$ case, with the difference that the approximate independent $\kappa$ values in RI are reached to higher temperatures and, therefore, there is a smaller window of $T_{\text{end}}$ to reproduce the current DM relic density. Finally, in Figure 10, obtained for $\omega=2/5$, there is no existence of RI since the values of $\kappa$ are greater than $1$. In other words, $\rho_{\phi,\text{ini}}$ is always greater than $\rho_{\gamma,\text{ini}}$. The parameters for NSC with and without bulk viscosity are very close (less than one order of magnitude) and, therefore, the FIMP DM candidates are not so sensitive to the inclusion of the bulk viscosity when the value of $\omega$ is higher compared these two cosmological scenarios. This also explains the absence of the red and blue zones, and the lines for $T_{\text{eq}}$, $T_{\text{c}}$, and $T_{\text{end}}$.
@@
III.2 Parameter spaces in the bulk viscous non-standard cosmologies
@@
From now on, we focus the analysis on the study of the parameter spaces ($T_{\text{end}}$,$\kappa$) of the model and $(m_{\chi},\langle\sigma v\rangle)$ of the FIMP DM candidate for the bulk viscous NSC scenario alone.
@@
In Figure 12, we depict the parameter space $(T_{\text{end}},\kappa)$ for $\hat{\xi_{0}}=10^{-2}$, $\langle\sigma v\rangle=10^{-22}$ GeV-2, and $\omega=0$. The solid blue, red, and green lines correspond to the parameter space that reproduces the current DM relic density for $m_{\chi}=10^{4}$, $10^{2}$, and $1$ GeV, respectively. The green and red zones represent the parameter space in which $\rho_{\phi}<\rho_{\gamma}$, $\forall t$, for $m_{\chi}=1$ and $10^{2}$ GeV, respectively. The grey zone corresponds to the forbidden BBN epoch that starts at $T_{\text{BBN}}\sim 4\times 10^{-3}$ GeV. From the figure, it can be observed that lower values of $m_{\chi}$ shift the curves downward to the left, while the regions in which $\rho_{\phi}<\rho_{\gamma}$ are shifted upward to the left. Hence, the blue zone is achieved at higher values of $T_{\text{end}}$ and lower values of $\kappa$. Interestingly, for $m_{\chi}=10^{4}$ GeV, exists a region in which the curve that reproduces the current DM relic density goes to higher values of $\kappa$. While this behavior occurs in the forbidden BBN zone, this parameter space may be accessible for higher values of the DM mass. Otherwise, Figure 13 shows the same parameter space as Figure 12 for $\hat{\xi_{0}}=10^{-2}$, $m_{\chi}=100$ GeV, and $\omega=0$. The solid blue, red, and green lines correspond to the parameter space that reproduces the current DM relic density for $\langle\sigma v\rangle=10^{-22}$, $10^{-24}$, and $10^{-26}$ GeV-2, respectively. The grey zones represent the parameter space in which $\rho_{\phi}<\rho_{\gamma}$, $\forall t$ (for the three cases), and the forbidden BBN epoch. As it can be noted, a decrease in the values for the thermally averaged DM production cross-section shifts the parameters that reproduce the current DM relic density downward to the right, allowing access to the “approximated independent $\kappa$-zone” described in Section III.1. Also, it is interesting to analyze the behavior of varying the DM mass (Figure 12) and its interactions (Figure 13). For instance, for higher values of $m_{\chi}$ appear a zone allowed to reproduce the current DM relic density in which the values of $\kappa$ increase when $T_{\text{end}}$ is decreasing. On the other hand, smaller values of $\langle\sigma v\rangle$ allow lower values of $\kappa$ when $T_{\text{end}}$ is diminishing up to the independent zone mentioned above. Finally, Figure 14 shows the parameter space mentioned before for $m_{\chi}=100$ GeV, $\langle\sigma v\rangle=10^{-22}$ GeV-2, and $\omega=0$. The solid green, red, blue and magenta lines correspond to the parameter space that reproduces the current DM relic density for $\hat{\xi}_{0}=10^{-3}$, $2.5\times 10^{-2}$, $5\times 10^{-2}$ and $10^{-1}$, respectively. The green, red, and blue zones represent their respective parameter space in which $\rho_{\phi}<\rho_{\gamma}$, $\forall t$. As before, the grey zone corresponds to the forbidden BBN epoch. As it is possible to see, higher values of $\hat{\xi_{0}}$ shift the curves and the regions, in which $\rho_{\phi}<\rho_{\gamma}$, downward to the right. Hence, the magenta zone is achieved at higher values of $T_{\text{end}}$ and lower values of $\kappa$. Also, this increment in the $\hat{\xi_{0}}$-values generates more prominent curvatures in the “approximately independent $\kappa$-zone”, reducing the $T_{\text{end}}$ parameter space that could reproduce the current DM relic density.
@@
III.3 FIMPs for a non-constant totally thermal averaged DM production cross-section
@@
IV Conclusions
In this paper, we revisit the bulk viscous NSC scenario studied in [106] in which the early universe is dominated by two interacting fluids, namely the new field $\phi$ and radiation, by considering that $\phi$ experiences dissipative processes during their cosmic evolution in the form of bulk viscosity. Working in the framework of Eckart’s theory for non-perfect fluids, we consider an interaction term of the form $\Gamma_{\phi}\rho_{\phi}$ and a bulk viscosity described by the expression $\xi=\xi_{0}\rho_{\phi}^{1/2}$. The latter has the advantage that when the field fully decays due to the interaction, the dissipation is negligible, and the standard $\Lambda$CDM cosmology is fully recovered. Also, when $\xi_{0}\to 0$, we recover the classical NSC scenario. Following the same scheme as in [106], we investigate the parameter space that can accurately reproduce the current DM relic density by adjusting both, the model and DM parameters within this novel NSC framework for FIMP DM candidates.
The main finding of our work is that the inclusion of bulk viscosity in NSCs brings to light DM parameters that are disregarded in both, $\Lambda$CDM and the classical NSC scenarios, as it was shown in Figure 7, giving us new windows to search for FIMP DM candidates. In particular, modifying the values of $m_{\chi}$ and $\langle\sigma v\rangle$ can shift the parameters ($T_{\text{end}}$, $\kappa$) that reproduce the current DM relic density. For example, an increment in the values of $m_{\chi}$ and $\langle\sigma v\rangle$ can shift the parameter space upward to the right (Figure 12) and upward to the left (Figure 13), respectively. These displacements can be combined to reach new windows of allowed parameters if a signal of DM is experimentally detected. In the same line, larger values of $\hat{\xi_{0}}$ (Figure 14) make more prominent the curvature to the approximately independent $\kappa$-zone (as it was described in Section III.1) and shift the parameters downward to the right. On the other hand, modifying the model parameters shifts the ($m_{\chi}$, $\langle\sigma v\rangle$) values that reproduce the current DM relic density. In particular, increasing the values of $\kappa$ and $T_{\text{end}}$ shift the parameter space upward (Figure 15) and downward to the right (Figure 16), respectively, reaching larger values of $\langle\sigma v\rangle$ and $m_{\chi}$. In this sense, larger values of $\hat{\xi_{0}}$ (Figure 17) make significant imprints in the DM parameters reaching larger values of $\langle\sigma v\rangle$.
@@ IV Conclusions
Finally, the case with non-constant $\langle\sigma v\rangle$ is studied for two scenarios: 1) for an operator of mass dimension 6 $(n=2)$ and 2) for an operator of mass dimension 7 $(n=4)$. In the first case, Figure 21 shows the same behavior previously observed for a constant value of $\langle\sigma v\rangle$ when the value of $\hat{\xi_{0}}$ changes. In contrast, Figure 22 shows that $\Lambda$ can take two distinct values for the same DM mass, a feature that does not arise in the case of a constant total thermally averaged DM production cross-section. Additionally, higher values of $\hat{\xi_{0}}$ shift the curves downward to the left. In the $n=4$ case (Figure 23), increasing $\hat{\xi}_{0}$ shifts the curves downward to the right.
Therefore, this paper is a further step in the study of FIMPs as DM candidates and a first approximation to highlight the imprints that the bulk viscosity can leave in these particles and their relic density in the early universe through a NSC scenario. In this sense, this work, in collaboration with Ref. [106], provides a comprehensive set of possibilities for opening new avenues of research in the study of FIMPs and WIMPs as DM candidates.
Acknowledgments
E.G. acknowledges the scientific support of Núcleo de Investigación No. 7 UCN-VRIDT 076/2020, Núcleo de Modelación y Simulación Científica (NMSC).

### v3

@@ 3.1 Comparison between scenarios
Figure 7: Comparison in the yield production between the classical NSC (solid red line) and the NSC with bulk viscosity (solid blue line) for $\kappa=10^{-3}$, $T_{\text{end}}=7\times 10^{-3}$ GeV, $\hat{\xi_{0}}=10^{-2}$, $m_{\chi}=100$ GeV, $\langle\sigma v\rangle=5\times 10^{-26}$ GeV-2, and $\omega=0$. The dashed and dotted lines correspond to $x_{\text{eq}}$ (cyan), $x_{\text{c}}$ (grey), and $x_{\text{end}}$ (green) for the NSC with and without bulk viscosity, respectively; while the dashed-dotted magenta line is the time when the DM candidate Freeze-In their number at $x_{\text{fi}}$. The green zone represents the current DM relic density according to Ref. Aghanim et al. (2020), observing that the parameter space considered in this case gives us the right amount of DM in the bulk viscous NSC scenario.
@@
The comparison in the parameter space of the NSC models, namely $\kappa$ and $T_{\text{end}}$, with a fixed $\omega$, is depicted in Figures 8, 9, and 10. The figures were obtained for the values $\hat{\xi_{0}}=10^{-2}$, $m_{\chi}=100$ GeV, and $\langle\sigma v\rangle=10^{-26}$ GeV-2. The solid red and blue lines correspond to the parameter space that reproduces the current DM relic density for the classical NSC and the NSC with bulk viscosity, respectively. We also delimit the three regions of interest (see Section 2.1) through the equalities $T_{\text{eq}}=T_{\text{fi}}$ (cyan), $T_{\text{c}}=T_{\text{fi}}$ (grey), and $T_{\text{end}}=T_{\text{fi}}$ (green) for the NSC with bulk viscosity (dashed lines) and the classical NSC (dashed-dotted lines). The red and blue zones represent their respective parameter space in which $\rho_{\phi}<\rho_{\gamma}$, $\forall t$, i.e., the regions in which the Hubble parameter is always approximately similar to the $\Lambda$CDM model ($H\sim T^{2}$). The grey zone corresponds to the forbidden BBN epoch that starts at $T_{\text{BBN}}\sim 4\times 10^{-3}$ GeV. In Figure 8, obtained for $\omega=0$, it can be seen that for higher values of $\kappa$ and $T_{\text{end}}$ (RIII), the parameter space is slightly the same, meanwhile in RII the curves separate from each other. However, the greatest differences between the NSC with and without bulk viscosity cases are presented in RI, exhibiting an approximate independent behavior of the $\kappa$ parameter, which can be explained by the dominance of the viscous term over the decaying term. On the other hand, Figure 9, obtained for $\omega=-2/5$, exhibits a similar behavior to the $\omega=0$ case, with the difference that the approximate independent $\kappa$ values in RI are reached to higher temperatures and, therefore, there is a smaller window of $T_{\text{end}}$ to reproduce the current DM relic density. Finally, in Figure 10, obtained for $\omega=2/5$, there is no existence of RI since the values of $\kappa$ are greater than $1$. In other words, $\rho_{\phi,\text{ini}}$ is always greater than $\rho_{\gamma,\text{ini}}$. The parameters for NSC with and without bulk viscosity are very close (less than one order of magnitude) and, therefore, the FIMP DM candidates are not so sensitive to the inclusion of the bulk viscosity when the value of $\omega$ is higher compared these two cosmological scenarios. This also explains the absence of the red and blue zones, and the lines for $T_{\text{eq}}$, $T_{\text{c}}$, and $T_{\text{end}}$.
@@
3.2 Parameter spaces in the bulk viscous non-standard cosmologies
@@
From now on, we focus the analysis on the study of the parameter spaces ($T_{\text{end}}$,$\kappa$) of the model and $(m_{\chi},\langle\sigma v\rangle)$ of the FIMP DM candidate for the bulk viscous NSC scenario alone. It is important to mention that, from this point onward, the choice of values for the free parameters of the model and for the DM is solely intended to illustrate the behavior of DM production within this new cosmological scenario and it is not a preference for an arbitrary value.
@@
In Figure 12, we depict the parameter space $(T_{\text{end}},\kappa)$ for $\hat{\xi_{0}}=10^{-2}$, $\langle\sigma v\rangle=10^{-22}$ GeV-2, and $\omega=0$. The solid blue, red, and green lines correspond to the parameter space that reproduces the current DM relic density for $m_{\chi}=10^{4}$, $10^{2}$, and $1$ GeV, respectively. The green and red zones represent the parameter space in which $\rho_{\phi}<\rho_{\gamma}$, $\forall t$, for $m_{\chi}=1$ and $10^{2}$ GeV, respectively. The grey zone corresponds to the forbidden BBN epoch that starts at $T_{\text{BBN}}\sim 4\times 10^{-3}$ GeV. From the figure, it can be observed that lower values of $m_{\chi}$ shift the curves downward to the left, while the regions in which $\rho_{\phi}<\rho_{\gamma}$ are shifted upward to the left. Hence, the blue zone is achieved at higher values of $T_{\text{end}}$ and lower values of $\kappa$. Interestingly, for $m_{\chi}=10^{4}$ GeV, exists a region in which the curve that reproduces the current DM relic density goes to higher values of $\kappa$. While this behavior occurs in the forbidden BBN zone, this parameter space may be accessible for higher values of the DM mass. Otherwise, Figure 13 shows the same parameter space as Figure 12 for $\hat{\xi_{0}}=10^{-2}$, $m_{\chi}=100$ GeV, and $\omega=0$. The solid blue, red, and green lines correspond to the parameter space that reproduces the current DM relic density for $\langle\sigma v\rangle=10^{-22}$, $10^{-24}$, and $10^{-26}$ GeV-2, respectively. The grey zones represent the parameter space in which $\rho_{\phi}<\rho_{\gamma}$, $\forall t$ (for the three cases), and the forbidden BBN epoch. As it can be noted, a decrease in the values for the thermally averaged DM production cross-section shifts the parameters that reproduce the current DM relic density downward to the right, allowing access to the “approximated independent $\kappa$-zone” described in Section 3.1. Also, it is interesting to analyze the behavior of varying the DM mass (Figure 12) and its interactions (Figure 13). For instance, for higher values of $m_{\chi}$ appear a zone allowed to reproduce the current DM relic density in which the values of $\kappa$ increase when $T_{\text{end}}$ is decreasing. On the other hand, smaller values of $\langle\sigma v\rangle$ allow lower values of $\kappa$ when $T_{\text{end}}$ is diminishing up to the independent zone mentioned above. Finally, Figure 14 shows the parameter space mentioned before for $m_{\chi}=100$ GeV, $\langle\sigma v\rangle=10^{-22}$ GeV-2, and $\omega=0$. The solid green, red, blue and magenta lines correspond to the parameter space that reproduces the current DM relic density for $\hat{\xi}_{0}=10^{-3}$, $2.5\times 10^{-2}$, $5\times 10^{-2}$ and $10^{-1}$, respectively. The green, red, and blue zones represent their respective parameter space in which $\rho_{\phi}<\rho_{\gamma}$, $\forall t$. As before, the grey zone corresponds to the forbidden BBN epoch. As it is possible to see, higher values of $\hat{\xi_{0}}$ shift the curves and the regions, in which $\rho_{\phi}<\rho_{\gamma}$, downward to the right. Hence, the magenta zone is achieved at higher values of $T_{\text{end}}$ and lower values of $\kappa$. Also, this increment in the $\hat{\xi_{0}}$-values generates more prominent curvatures in the “approximately independent $\kappa$-zone", reducing the $T_{\text{end}}$ parameter space that could reproduce the current DM relic density.
@@
3.3 FIMPs for a non-constant totally thermal averaged DM production cross-section
@@
Larger couplings would typically be excluded in the standard $\Lambda$CDM scenario, but remain viable within our viscous NSC framework can also enhance the rate of decay processes. If the DM candidates are not absolutely stable but have lifetimes exceeding the age of the Universe, they may decay into SM particles such as photons or neutrinos. This opens the possibility for indirect detection, provided the signal is sufficiently strong, through gamma-ray lines or, in models involving neutrino final states, via detectors such as IceCube or Super-Kamiokande Barranco et al. (2018); Lambiase et al. (2018).
4 Conclusions
In this paper, we revisit the bulk viscous NSC scenario studied in González et al. (2024) in which the early universe is dominated by two interacting fluids, namely the new field $\phi$ and radiation, by considering that $\phi$ experiences dissipative processes during their cosmic evolution in the form of bulk viscosity. Working in the framework of Eckart’s theory for non-perfect fluids, we consider an interaction term of the form $\Gamma_{\phi}\rho_{\phi}$ and a bulk viscosity described by the expression $\xi=\xi_{0}\rho_{\phi}^{1/2}$. The latter has the advantage that when the field fully decays due to the interaction, the dissipation is negligible, and the standard $\Lambda$CDM cosmology is fully recovered. Also, when $\xi_{0}\to 0$, we recover the classical NSC scenario. Following the same scheme as in González et al. (2024), we investigate the parameter space that can accurately reproduce the current DM relic density by adjusting both, the model and DM parameters within this novel NSC framework for FIMP DM candidates.
The main finding of our work is that the inclusion of bulk viscosity in NSCs brings to light DM parameters that are disregarded in both, $\Lambda$CDM and the classical NSC scenarios, as it was shown in Figure 7, giving us new windows to search for FIMP DM candidates. In particular, modifying the values of $m_{\chi}$ and $\langle\sigma v\rangle$ can shift the parameters ($T_{\text{end}}$, $\kappa$) that reproduce the current DM relic density. For example, an increment in the values of $m_{\chi}$ and $\langle\sigma v\rangle$ can shift the parameter space upward to the right (Figure 12) and upward to the left (Figure 13), respectively. These displacements can be combined to reach new windows of allowed parameters if a signal of DM is experimentally detected. In the same line, larger values of $\hat{\xi_{0}}$ (Figure 14) make more prominent the curvature to the approximately independent $\kappa$-zone (as it was described in Section 3.1) and shift the parameters downward to the right. On the other hand, modifying the model parameters shifts the ($m_{\chi}$, $\langle\sigma v\rangle$) values that reproduce the current DM relic density. In particular, increasing the values of $\kappa$ and $T_{\text{end}}$ shift the parameter space upward (Figure 15) and downward to the right (Figure 16), respectively, reaching larger values of $\langle\sigma v\rangle$ and $m_{\chi}$. In this sense, larger values of $\hat{\xi_{0}}$ (Figure 17) make significant imprints in the DM parameters reaching larger values of $\langle\sigma v\rangle$.
@@ 4 Conclusions
Finally, the case with non-constant $\langle\sigma v\rangle$ is studied for two scenarios: 1) for an operator of mass dimension 6 $(n=2)$ and 2) for an operator of mass dimension 7 $(n=4)$. In the first case, Figure 21 shows the same behavior previously observed for a constant value of $\langle\sigma v\rangle$ when the value of $\hat{\xi_{0}}$ changes. In contrast, Figure 22 shows that $\Lambda$ can take two distinct values for the same DM mass, a feature that does not arise in the case of a constant total thermally averaged DM production cross-section. Additionally, higher values of $\hat{\xi_{0}}$ shift the curves downward to the left. In the $n=4$ case (Figure 23), increasing $\hat{\xi}_{0}$ shifts the curves downward to the right. The larger couplings that are needed to set the current relic density opens the possibility for indirect detection in next-generation experiments.
Therefore, this paper is a further step in the study of FIMPs as DM candidates and a first approximation to highlight the imprints that the bulk viscosity can leave in these particles and their relic density in the early universe through a NSC scenario. In this sense, this work, in collaboration with Ref. González et al. (2024), provides a comprehensive set of possibilities for opening new avenues of research in the study of FIMPs and WIMPs as DM candidates. On the other hand, including various cosmological observations, such as the CMB, LSS, or BBN, would allow for tighter constraints on the dark matter parameters. This will be explored in future work.
[…]
Conceptualization, E.G., C.M., N.S.M. and R.S.; methodology, E.G., C.M., N.S.M. and R.S.; software, C.M., N.S.M. and R.S.; validation, E.G. and C.M.; formal analysis, E.G., C.M., N.S.M. and R.S.; investigation, E.G., C.M., N.S.M. and R.S.; data curation, C.M., N.S.M. and R.S.; writing—original draft preparation, E.G. and C.M.; writing—review and editing, E.G., C.M., N.S.M. and R.S.; visualization, E.G., C.M., N.S.M. and R.S.; supervision, E.G. and C.M.; project administration, E.G. and C.M. All authors have read and agreed to the published version of the manuscript.
[…]
This research received no external funding.
[…]
No new data were generated during the development of this work.
Acknowledgements.
[…]
DM Dark Matter
DE Dark Energy
CDM Cold Dark Matter
WIMPs Weakly Interacting Massive Particles
FIMPs Feebly Interacting Massive Particles
SM tandard Model
WDM Warm Dark Matter
NSCs Non-Standard Cosmologies
BBN Big Bang Nucleosynthesis
RI Region I
RII Region II
RIII Region III
RVI Region VI
UV UltraViolet
IR InfraRed
[…]